Prev Next

SW&IT Meeting Minutes 2018-12-24

Attendees: Andrew, David, Igor, Matthieu, Petr, Stephen, Xianyu, Yang


Labor Reporting

Report auto-generated every, Do not edit this table, use entries in smartsheets.

LR20181203 - Bec, Matthieu 65.0
LR20181203 - Colson, Andrew 48.0
LR20181203 - Edmeades, David 64.0
LR20181203 - Sola, Igor 64.0
LR20181203 - Kubanek, Petr 80.0
LR20181203 - Hooper, Stephen 80.0
LR20181203 - Zhang, Yang 80.0
LR20181203 - Zhang, Xianyu 27.0
LR20181217 - Bec, Matthieu 80.0
LR20181217 - Colson, Andrew 32.0
LR20181217 - Edmeades, David 40.0
LR20181217 - Sola, Igor 40.0
LR20181217 - Kubanek, Petr 40.0
LR20181217 - Hooper, Stephen 40.0
LR20181217 - Zhang, Yang 40.0
LR20181217 - Zhang, Xianyu 40.0

Transition plan

Handover steps

  1. primary + secondary do self learning, gather information, read wiki
  2. primary organize teamviewer/zoom walk through with Kellee and secondary
  3. primary draft techtalk w. input from secondary
  4. techtalk review by Kellee
  5. primary does a team presentation

Please update this table to indicate progress.

Application(s)Sorted ascending Primary Secondary Week
ALH Igor(2)
Matthieu (2)
AC (2)
Petr (2)
AC (2)
Matthieu (2)
INDI Igor (2)
Matthieu (2)
Yang (5)
Igor (1/2) 2
particle monitors AC Petr 4
Status Server Stephen (2)
Matthieu (2)
Yang (1)
Stephen (1) 3
AC (2)
Petr (2)

News from management

  • PEPSI/pol high profile run in January: AGW3 & 4 must be ready
  • ARGOS half-night Dec 22 - confirmed
  • Time Sheets for Pay Period End Date December 30, 2018 must be submitted and approved no later than 5:00 pm on Tuesday, December 18, 201

Previous Action Items

  • Skill set list (DE)
  • UA time (AC)

Open topics

Bring any topics you want for discussion at the meeting here. Use <topic>[<name>] formatting.

  • seeing monitor upgrade
  • end of year support
  • network dropout - antenna monitoring?
  • iLocater Single-Mode Data Fibers - CRA/CRB to Electronics Shop
  • BCU f/w upgrade documentation
  • Dust monitor status
  • AO improvement: Luci kwds
  • AO cabinet switch upgrade - follow up
  • wifi summit upgrade (note: end-of-year deadline)
  • Tucson CR power overhaul plan (note: end-of-year deadline)
  • n507/obs upgrade (note: end-of-year deadline)
  • synology new drives
  • PGray mirror cover req. feedback
  • Installing python packages on flao-sxwfs
  • support aligned with mountain schedule (starting Tuesday instead of Monday)? (PK)

This week in github

Team reports


  • Finished the intro MODS wiki page.
  • Completed the MODS tech talk.
  • A preliminary web interface for AO-Cravat is constructed.
  • Meeting with Al to discuss progress and plans for the AO-Cravat project.
  • Next will focus on developing an infrastructure for server part, will start with reading and experimenting with mod_python and Apache settings.


  • New Synology drives arrived finally
    • Installed (no immediate problem with the 12TB drives, despite not being explicitely approved by Synology)
    • Arrays and volumes building
    • By the time anyone reads this, they should be available for use
      • Volume2: 43.64TiB
      • Volume4: 32.73TiB
      • Both RAID-6
      • Removed the spanning of drives between units; now all volumes are fully contained within one hardware container.
  • Replaced batteries in Rack1 UPS
  • Attempt to modernize AGw computer OS software to read modern USB storage devices, but this was unsuccessful. Suspect a limit of the Windows 7 driver.
    • As an alternate, we connected a Samba share on the OAC machine which is visible to AGw and LBT network clients
  • Autodesk software meeting; don't think our internal plans have changed re: 2018 everything


  • This Week
    • LUCI
      • Added AFC visualization to obs# production (test) directory for Dave T. to try.
      • Improve on Auxiliary cabinet exchange step. Basically use a retry if a switch gets stuck. A version of the code is in production but not tested. It might be tested in the next exchange.
      • Fixed AOPMDEC Fits keyword
      • Looking at RTD issue #7473
        • Unable to use AFC Align Pinhole
      • Added new pinhole definitions to LUCI 1 tables.
  • Next Week (January)
    • Possibly working on OT ARGOS/ESM binocular scripting
    • Look at RTD issues #7473, #7488 and #7512.
      • Unable to use AFC Align Pinholes to Align Field Stop For N30 Camera
      • Slow subtraction in RTD
      • LUCI-2 skips field image during MOS acquisition




  • Microwave web camera changes.
  • analysis host reinstallation as Fedora 29 for Matthieu.
  • moved filesystems on Synology for new disk installation, and requisite changes on mounting hosts.
  • IPC kickstart install for OVMS-TT
  • Re-quoted ARGOS computer, and asked that Matthieu be added to Dell Premier.
  • UITS complaints that AO was DOS-ing a computer in the Netherlands.
  • Powered off old AO machines
  • Powered off 32-bit


last week:
  • Generate the excel memory map for old and new crate BCU memory map
  • Analysis thousands measurements for current 4D resolution check
  • OVMS+TT meeting
  • Fastlink fiber swaps configration and debug
  • Rewritten/Reformat the SW requirements with Matthieu
  • Change mis-typed LN IP on AdSec machine because UITS complains
  • Dicussion about SX AdSec TT slow issue
  • Check "special" housekeeper parameters


  • OC, WPC, Management
  • OVMS+TT meeting, received 3xIPC, received cable serial specs passed on to Eng. for procurement
  • ao-cravat meeting and way forward, meet once a month at least
  • Ketiv support call, Bucki moving on
  • attend LUCI remote test from Italy using vnc
  • refactoring and packaging AHS (alarm handling system)
  • schedule 3x phone screening interview
  • schedule sci-arch meeting mid Jan,
  • update my skillset, send to CV
  • misc github activities
  • AO keyword planning with Andrew and Igor
  • deltatau site training, waiting clean invoice
  • review 4D s/w requirements
  • engineering science archive demo on analysis
  • site training with deltatau, get invoice
  • phone screening
Topic revision: r9 - 31 Dec 2018, MatthieuBec
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback